# Transaction Fee Market Design for Parallel Execution

### v1

@@ 3. GCM Properties
  • Transactions with overlapping resource access sets cannot be executed in parallel: one must finish before the other can begin.
timer1r2r3r4r5r6r7r8resources
timer1r2r3r4r5r6r7r8resources
Figure 6. Illustration of optimal schedules for a set of four transactions: $\textit{tx}_{1}\simeq(2,\{r_{2},\dots,r_{8}\})$ (shown in   ), $\textit{tx}_{2}\simeq(4,\{r_{2},r_{3}\})$ (shown in   ), $\textit{tx}_{3}\simeq(5,\{r_{4},r_{5},r_{6}\})$ (shown in   ), and $\textit{tx}_{4}\simeq(2,\{r_{7},r_{8}\})$ (shown in   ). In the left plot, we show the optimal schedule for $n=3$, and in the right plot for $n=2$. Notice how for $n=2$, we cannot schedule the three transactions $\textit{tx}_{2}$, $\textit{tx}_{3}$, and $\textit{tx}_{4}$ to be executed in parallel even though they access pairwise-disjoint sets of resources.
Note that some of these conditions are natural for scheduling in general, while others arise from us assuming the simple lock-based execution policy. [^7] Using locks is one way to enforce this policy, but in our case — where the contents of a block are known before execution commences — it can also be achieved without locks, as long as the execution environment ensures that threads strictly follow a pre-determined schedule. We chose this name because it is largely suggestive of the intended semantics. For our purposes, we are not concerned with which thread executes which transaction, but only that no more than $n$ transactions ever execute simultaneously (so we can draw concurrent schedules as in Fig. 6, which illustrates the concept). To determine gas consumption, our notation will need to capture even less — given a scheduler $s$ (for $n$ threads), we write $v^{s}(T)$ for the makespan of the schedule produced by $s$ for the set of transactions $T$ (i.e., the time required to execute the schedule in parallel using $n$ threads). It is instructive to read the paper having in mind as a prime example $s$ being the optimal scheduler (for $n$ threads), which returns a schedule for $n$ threads that minimizes the makespan. However, implementing such a scheduler is computationally prohibitive in practice, so greedy heuristics are typically used instead. To make our results apply even to non-optimal schedulers, we assume a set of minimal, reasonable properties for the scheduler:
@@
    Monotonicity in $t$ and $R$: consider any set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $R_{1}\subseteq R_{2}$, i.e., $\textit{tx}_{1}\lesssim\textit{tx}_{2}$, then we have $v(T\cup\{\textit{tx}_{1}\})\leq v(T\cup\{\textit{tx}_{2}\})$. Intuitively, “larger” transactions are no easier to schedule.
@@
An attentive reader may notice that (S1)(S3) partially resemble Properties 1, 2, 3, 4 and 5 for GCMs. This resemblance is not coincidental, but it is important to emphasize that they address different aspects: the former are properties of the scheduler, while the latter are properties of the GCM. A GCM can, in fact, satisfy Properties 1, 2, 3, 4 and 5 without depending on the scheduler at all (e.g., the current mechanism). Conversely, for mechanisms defined in terms of the scheduler, (S1)(S3) play a crucial role in establishing their properties, including Properties 1, 2, 3, 4 and 5.
@@
Property 6 (Scheduling Sensitivity).
@@
Intuitively, transactions with higher marginal contributions to the execution time should consume no less gas. [^10] Technically, what we wrote above in Property 6 are not marginal contributions, but can be made to be by subtracting $v(T)$ from both sides of the inequality in the antecedent. As usual, we define a strict version of this property, where the latter inequality also becomes strict (i.e., higher marginal contributions in execution time imply higher gas consumption). One might be tempted to believe that Scheduling Sensitivity implies Resource-Time Sensitivity. However, proving this requires a non-strict inequality in the premise $v(T\cup\{\textit{tx}_{1}\})<v(T\cup\{\textit{tx}_{2}\})$.
There is a second way in which gas consumption should adequately indicate the effort required to execute transactions (Property 7). Specifically, the gas consumption of all transactions in a block should collectively account for the total time needed to execute the block. This can be seen as properly tracking the time consumed by the execution environment to execute the block.
@@
Next, we introduce two practical properties. The first requires that transaction submitters be able to estimate a transaction’s gas consumption in advance. We formalize this by requiring that $\textsf{gas}_{T}(\textit{tx})$ does not depend on $T$ (Property 8).
@@
This property ensures a good user experience by making it easy for users to estimate gas usage. If the gas consumption of a transaction depended on the remaining transactions in the block, this estimation would become significantly more complex. In general, we aim to keep gas estimation straightforward to avoid giving an advantage to more sophisticated users. Additionally, as we will see later, a GCM satisfying this property can be seamlessly composed with existing TFMs, retaining their desirable properties (see Section 6). Sadly, as one might have already guessed, Easy Gas Estimation is incompatible with Scheduling Sensitivity (except for constant mechanisms, i.e., GCMs that return a constant independent of $T$ and tx), and also incompatible with Efficiency:
@@
Easy Gas Estimation (Property 8) is:
  1. (1)
    Incompatible with Scheduling Sensitivity (Property 6) unless using a constant GCM. [^11] Constant GCMs satisfy Scheduling Sensitivity, but not Strict Scheduling Sensitivity.
  2. (2)
    Incompatible with Efficiency (Property 7).
Proof.
@@
  1. (1)
    Assume that $M$ has Scheduling Sensitivity; we will show that $M$ is a constant mechanism.
    Call two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$ incomparable if neither $R_{1}\subseteq R_{2}$ nor $R_{2}\subseteq R_{1}$. As a first step, we will show that if $\textit{tx}_{1}$ and $\textit{tx}_{2}$ are incomparable, then $\textsf{gas}^{M}(\textit{tx}_{1})=\textsf{gas}^{M}(\textit{tx}_{2})$. To show this, assume that $R_{1}\not\subseteq R_{2}$. We will show that then $\textsf{gas}^{M}(\textit{tx}_{1})\geq\textsf{gas}^{M}(\textit{tx}_{2})$. Exchanging the roles of $\textit{tx}_{1}$ and $\textit{tx}_{2}$ will then give the conclusion. Let $r\in R_{1}\setminus R_{2}$ and $\textit{tx}_{3}$ be another transaction such that $\textit{tx}_{3}\simeq(t_{3},\{r\})$ for some $t_{3}>t_{2}$, and define $T=\{\textit{tx}_{3}\}$. Then, for any number of threads $n\geq 2$ and any scheduler that is optimal for two transactions, we have $v(T\cup\{\textit{tx}_{1}\})=t_{1}+t_{3}>t_{3}=v(T\cup\{\textit{tx}_{2}\})$. By Scheduling Sensitivity, this implies that $\textsf{gas}^{M}(\textit{tx}_{1})\geq\textsf{gas}^{M}(\textit{tx}_{2})$.
    We now know that $M$ associates the same gas consumption to any pair of incomparable transactions. Armed as such, consider two arbitrary transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$. Let $r$ be an arbitrary resource not in $R_{1}\cup R_{2}$ and $\textit{tx}_{3}$ be a transaction such that $\textit{tx}_{3}\simeq(1,\{r\})$. Since $R_{1}$ and $R_{2}$ are non-empty, one can easily see that $\textit{tx}_{3}$ is incomparable with both $\textit{tx}_{1}$ and $\textit{tx}_{2}$, from which $\textsf{gas}^{M}(\textit{tx}_{1})=\textsf{gas}^{M}(\textit{tx}_{3})=\textsf{gas}^{M}(\textit{tx}_{2})$.
  2. (2)
    Assume for a contradiction that $M$ has Efficiency, then for any transaction $\textit{tx}\simeq(t,R)$ we have $\textsf{gas}^{M}(\textit{tx})=\textsf{gas}_{\varnothing}^{M}(\{\textit{tx}\})=v(\{\textit{tx}\})=t$. Hence, by definition, $M$ is the current mechanism, which can be easily seen to not satisfy Efficiency: assume $n\geq 2$ threads and consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2}\}$ where $\textit{tx}_{1}\simeq(1,\{r_{1}\})$ and $\textit{tx}_{2}\simeq(1,\{r_{2}\})$. In this case, $\textsf{gas}^{M}(T)=\textsf{gas}^{M}(\textit{tx}_{1})+\textsf{gas}^{M}(\textit{tx}_{1})=1+1=2\neq 1=v(T)$. ∎
Our second practical property requires that gas consumption be efficiently computable, i.e., in polynomial time (Property 9). A GCM that does not satisfy this property would be unsuitable for the blockchain context, where gas computation is intended to be a straightforward component.
@@
4. GCM Proposals
We now propose multiple GCM designs. [^12] Note again that we will solely consider deterministic mechanisms. This is common for blockchain fee markets given the difficulty of having a true source of randomness on blockchains. Motivated by the incompatibilities in Theorem 3.4, these fall into two categories:
@@ 4. GCM Proposals
    Mechanisms with Easy Gas Estimation, in which each transaction’s gas consumption is computed in isolation. Such mechanisms tend to be both straightforward and attractive but can achieve neither Scheduling Sensitivity [^13] Except for constant GCMs. nor Efficiency.
@@
    Mechanisms without Easy Gas Estimation, which, given a set of transactions $T$, rely on $v(T)$ or, more generally, $\left(v(T^{\prime})\right)_{T^{\prime}\subseteq T^{\prime}}$ to holistically calculate gas consumptions for the block $T$, requiring knowledge of the entire block. Instead, these mechanisms aim to achieve Scheduling Sensitivity and/or Efficiency.
4.1. Mechanisms with Easy Gas Estimation
We already introduced the current GCM (Definition 2.3). For the next mechanism, we assume a globally available vector of positive weights for the resources $(w_{r})_{r\in\mathcal{R}}$. For instance, these weights could all be 1. Alternatively, higher-weight resources could correspond to resources under higher (historical) demand. For our purposes, we only need to assume that the weights are known and do not depend on the block being built. In practice, one could update the weights between blocks to accurately reflect resource demand (the exact details are not relevant here; see Section 6 for further discussion). Given the weights, we define:
@@ 4.1. Mechanisms with Easy Gas Estimation
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$ with $\textit{tx}\simeq(t,R)$, the Weighted Area GCM computes the amount of gas used by tx as follows:
(1) $\textsf{gas}_{T}^{\textit{WA}}(\textit{tx}):=t\cdot\left(1+\sum_{r\in R}w_{r}\right)$
@@
To gain intuition, it is instructive to consider the former case, where all weights are 1, i.e., the unweighted area mechanism, in which case Eq. 1 becomes $\textsf{gas}_{T}^{\textit{WA}}(\textit{tx})=t\cdot\left(1+|R|\right)$. This mechanism can also be viewed as the addition of two terms: the current term, i.e., the gas consumption of tx under the current GCM, namely $t$, and the area term, i.e., the area of a $t\times|R|$ rectangle (which is also how we draw transactions in our diagrams). The current term is helpful to ensure no transaction ever costs too little when the weights of all resources accessed by a transaction are too close to zero. [^14] This term has a practical motivation. In reality, transactions will be scheduled on a small finite number of threads. Thus, even if they do not access any high-demand resource, they occupy space in the schedule proportional to their execution time. Note, moreover, that this term could technically be omitted, and our results would still hold.
The area term, on the other hand, which is the main component, intuitively corresponds to “negated throughput.” That is, executing the transaction requires holding $|R|$ locks for $t$ time units, during which no other transactions requiring any of those resources can execute. In Fig. 7, we illustrate this idea. Transaction tx prevents the execution of other transactions across its entire resource set but only utilizes each resource for a short period. While the current GCM charges a transaction based solely on its total computation time (i.e., the height of the rectangle), the weighted area GCM also accounts for the resources for which it “negated throughput”, i.e., the area occupied by the transaction. This occupied area prevents other transactions using the same resources from being scheduled in parallel.
time6r1r2r3r4r5r6r7r8resources
Figure 7. Illustration of a worst-case transaction in terms of parallelizability. Transaction $\textit{tx}\simeq(6,\{r_{2},\dots,r_{7}\})$ (shown in   ) takes 6 units of time to execute and accesses 6 resources. However, the transaction only operates on each resource for one unit of time each (shown in   ).
Along the same line of reasoning, to further reinforce why this mechanism is a reasonable choice, consider a transaction set $T$. The sum $\sum_{\textit{tx}\in T}t(\textit{tx})\cdot|R(\textit{tx})|$ represents the “total area” of transactions in $T$. Dividing this term by the number of resources used in $T$ provides a lower bound on $v(T)$, serving as a rough proxy that does not require knowledge of $T$ when computing individual gas consumptions. Finally, we note that, even in the case of non-unit weights, the term weighted area remains meaningful, as the area term still corresponds to the area of the respective rectangle in our diagrams if we give the column of each resource $r\in\mathcal{R}$ a width of $w_{r}$.
4.2. Mechanisms without Easy Gas Estimation

### v2

@@ 3 GCM Properties
  • Transactions with overlapping storage key access sets cannot be executed in parallel: one must finish before the other can begin.
[…]
Figure 6: Illustration of optimal schedules for a set of four transactions: $\textit{tx}_{1}\simeq(2,\{k_{2},\dots,k_{8}\})$ (shown in   ), $\textit{tx}_{2}\simeq(4,\{k_{2},k_{3}\})$ (shown in   ), $\textit{tx}_{3}\simeq(5,\{k_{4},k_{5},k_{6}\})$ (shown in   ), and $\textit{tx}_{4}\simeq(2,\{k_{7},k_{8}\})$ (shown in   ). In the left plot, we show the optimal schedule for $n=3$, and in the right plot for $n=2$. Notice how for $n=2$, we cannot schedule the three transactions $\textit{tx}_{2}$, $\textit{tx}_{3}$, and $\textit{tx}_{4}$ to be executed in parallel even though they access pairwise-disjoint sets of storage keys.
Note that some of these conditions are natural for scheduling in general, while others arise from us assuming the simple lock-based execution policy. [^10] Using locks is one way to enforce this policy, but in our case — where the contents of a block are known before execution commences — it can also be achieved without locks, as long as the execution environment ensures that threads strictly follow a pre-determined schedule. We chose this name because it is largely suggestive of the intended semantics. For our purposes, we are not concerned with which thread executes which transaction, but only that no more than $n$ transactions ever execute simultaneously (so we can draw concurrent schedules as in Figure˜6, which illustrates the concept). To determine gas consumption, our notation will need to capture even less — given a scheduler $s$ (for $n$ threads), we write $v^{s}(T)$ for the makespan of the schedule produced by $s$ for the set of transactions $T$ (i.e., the time required to execute the schedule in parallel using $n$ threads). It is instructive to read the paper having in mind as a prime example $s$ being the optimal scheduler (for $n$ threads), which returns a schedule for $n$ threads that minimizes the makespan. However, implementing such a scheduler is computationally prohibitive in practice, so greedy heuristics are typically used instead. To make our results apply even to non-optimal schedulers, we assume a set of minimal, reasonable properties for the scheduler:
@@
    Monotonicity in $t$ and $K$: consider any set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$, i.e., $\textit{tx}_{1}\lesssim\textit{tx}_{2}$, then we have $v(T\cup\{\textit{tx}_{1}\})\leq v(T\cup\{\textit{tx}_{2}\})$. Intuitively, “larger” transactions are no easier to schedule.
@@
An attentive reader may notice that (S1)(S3) partially resemble Properties˜1, 2, 3, 4 and 5 for GCMs. This resemblance is not coincidental, but it is important to emphasize that they address different aspects: the former are properties of the scheduler, while the latter are properties of the GCM. A GCM can, in fact, satisfy Properties˜1, 2, 3, 4 and 5 without depending on the scheduler at all (e.g., the current mechanism). Conversely, for mechanisms defined in terms of the scheduler, (S1)(S3) play a crucial role in establishing their properties, including Properties˜1, 2, 3, 4 and 5.
@@
Property 6 (Scheduling Monotonicity).
@@
Intuitively, transactions with higher marginal contributions to the execution time should consume no less gas. [^13] Technically, what we wrote above in Property 6 are not marginal contributions, but can be made to be by subtracting $v(T)$ from both sides of the inequality in the antecedent. As usual, we define a strict version of this property, where the latter inequality also becomes strict (i.e., higher marginal contributions in execution time imply higher gas consumption). One might be tempted to believe that Scheduling Monotonicity implies Storage Key-Time Monotonicity. However, proving this requires a non-strict inequality in the premise $v(T\cup\{\textit{tx}_{1}\})<v(T\cup\{\textit{tx}_{2}\})$.
There is a second way in which gas consumption should adequately indicate the effort required to execute transactions (Property˜7). Specifically, the gas consumption of all transactions in a block should collectively account for the total time needed to execute the block. This can be seen as properly tracking the time consumed by the execution environment to execute the block.
@@
Next, we introduce two practical properties. The first requires that transaction submitters be able to estimate a transaction’s gas consumption in advance. We formalize this by requiring that $\textsf{gas}_{T}(\textit{tx})$ does not depend on $T$ (Property˜8).
@@
This property ensures a good user experience by making it easy for users to estimate gas usage. If the gas consumption of a transaction depended on the remaining transactions in the block, this estimation would become significantly more complex. In general, we aim to keep gas estimation straightforward to avoid giving an advantage to more sophisticated users. Additionally, as we will see later, a GCM satisfying this property can be seamlessly composed with existing TFMs, retaining their desirable properties (see Section˜6). Sadly, as one might have already guessed, Easy Gas Estimation is incompatible with Scheduling Monotonicity (except for constant mechanisms, i.e., GCMs that return a constant independent of $T$ and tx), and also incompatible with Efficiency:
@@
Easy Gas Estimation (Property˜8) is:
  1. 1.
    Incompatible with Scheduling Monotonicity (Property˜6) unless using a constant GCM. [^14] Constant GCMs satisfy Scheduling Monotonicity, but not Strict Scheduling Monotonicity.
  2. 2.
    Incompatible with Efficiency (Property˜7).
Proof 3.5.
@@
  1. 1.
    Assume that $M$ has Scheduling Monotonicity; we will show that $M$ is a constant mechanism.
    Call two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$ incomparable if neither $K_{1}\subseteq K_{2}$ nor $K_{2}\subseteq K_{1}$. As a first step, we will show that if $\textit{tx}_{1}$ and $\textit{tx}_{2}$ are incomparable, then $\textsf{gas}^{M}(\textit{tx}_{1})=\textsf{gas}^{M}(\textit{tx}_{2})$. To show this, assume that $K_{1}\not\subseteq K_{2}$. We will show that then $\textsf{gas}^{M}(\textit{tx}_{1})\geq\textsf{gas}^{M}(\textit{tx}_{2})$. Exchanging the roles of $\textit{tx}_{1}$ and $\textit{tx}_{2}$ will then give the conclusion. Let $k\in K_{1}\setminus K_{2}$ and $\textit{tx}_{3}$ be another transaction such that $\textit{tx}_{3}\simeq(t_{3},\{k\})$ for some $t_{3}>t_{2}$, and define $T=\{\textit{tx}_{3}\}$. Then, for any number of threads $n\geq 2$ and any scheduler that is optimal for two transactions, we have $v(T\cup\{\textit{tx}_{1}\})=t_{1}+t_{3}>t_{3}=v(T\cup\{\textit{tx}_{2}\})$. By Scheduling Monotonicity, this implies that $\textsf{gas}^{M}(\textit{tx}_{1})\geq\textsf{gas}^{M}(\textit{tx}_{2})$.
    We now know that $M$ associates the same gas consumption to any pair of incomparable transactions. Armed as such, consider two arbitrary transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$. Let $k$ be an arbitrary storage key not in $K_{1}\cup K_{2}$ and $\textit{tx}_{3}$ be a transaction such that $\textit{tx}_{3}\simeq(1,\{r\})$. Since $K_{1}$ and $K_{2}$ are non-empty, one can easily see that $\textit{tx}_{3}$ is incomparable with both $\textit{tx}_{1}$ and $\textit{tx}_{2}$, from which $\textsf{gas}^{M}(\textit{tx}_{1})=\textsf{gas}^{M}(\textit{tx}_{3})=\textsf{gas}^{M}(\textit{tx}_{2})$.
  2. 2.
    Assume for a contradiction that $M$ has Efficiency, then for any transaction $\textit{tx}\simeq(t,K)$ we have $\textsf{gas}^{M}(\textit{tx})=\textsf{gas}_{\varnothing}^{M}(\{\textit{tx}\})=v(\{\textit{tx}\})=t$. Hence, by definition, $M$ is the current mechanism, which can be easily seen to not satisfy Efficiency: assume $n\geq 2$ threads and consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$ and $\textit{tx}_{2}\simeq(1,\{k_{2}\})$. In this case, $\textsf{gas}^{M}(T)=\textsf{gas}^{M}(\textit{tx}_{1})+\textsf{gas}^{M}(\textit{tx}_{1})=1+1=2\neq 1=v(T)$.
Our second practical property requires that gas consumption be efficiently computable, i.e., in polynomial time (Property˜9). A GCM that does not satisfy this property would be unsuitable for the blockchain context, where gas computation is intended to be a straightforward component.
@@
4 GCM Proposals
We now propose multiple GCM designs. Motivated by the incompatibilities in Theorem˜3.4, these fall into two categories:
@@ 4 GCM Proposals
    Mechanisms with Easy Gas Estimation, in which each transaction’s gas consumption is computed in isolation. Such mechanisms tend to be both straightforward and attractive but can achieve neither Scheduling Monotonicity [^15] Except for constant GCMs. nor Efficiency.
@@
    Mechanisms without Easy Gas Estimation, which, given a set of transactions $T$, rely on $v(T)$ or, more generally, $\left(v(T^{\prime})\right)_{T^{\prime}\subseteq T^{\prime}}$ to holistically calculate gas consumptions for the block $T$, requiring knowledge of the entire block. Instead, these mechanisms aim to achieve Scheduling Monotonicity and/or Efficiency.
4.1 Mechanisms with Easy Gas Estimation
We already introduced the current GCM (Definition˜2.3). For the next mechanism, we assume a globally available vector of positive weights for the storage keys $(w_{k})_{k\in\mathcal{K}}$. For instance, these weights could all be 1. Alternatively, higher-weight storage keys could correspond to storage keys under higher (historical) demand. For our purposes, we only need to assume that the weights are known and do not depend on the block being built. In practice, one could update the weights between blocks to accurately reflect storage key demand (the exact details are not relevant here; see Section˜6 for further discussion). Given the weights, we define:
@@ 4.1 Mechanisms with Easy Gas Estimation
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$ with $\textit{tx}\simeq(t,K)$, the Weighted Area GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{\textit{WA}}(\textit{tx}):=t\cdot\left(1+\sum_{k\in K}w_{k}\right)$ (1)
@@
To gain intuition, it is instructive to consider the former case, where all weights are 1, i.e., the unweighted area mechanism, in which case Equation˜1 becomes $\textsf{gas}_{T}^{\textit{WA}}(\textit{tx})=t\cdot\left(1+|K|\right)$. This mechanism can also be viewed as the addition of two terms: the current term, i.e., the gas consumption of tx under the current GCM, namely $t$, and the area term, i.e., the area of a $t\times|K|$ rectangle (which is also how we draw transactions in our diagrams). The current term is helpful to ensure no transaction ever costs too little when the weights of all storage keys accessed by a transaction are too close to zero. [^16] This term has a practical motivation. In reality, transactions will be scheduled on a small finite number of threads. Thus, even if they do not access any high-demand storage key, they occupy space in the schedule proportional to their execution time. Note, moreover, that this term could technically be omitted, and our results would still hold.
[…]
4.2 Mechanisms without Easy Gas Estimation